OpenStudy (anonymous):

2sin^2 x + 3cosx = 3 Chapter uses trig functions

mathslover (mathslover):

You have to prove that or solve?

OpenStudy (anonymous):

start be recalling that \(\sin^2(x)=1-\cos^2(x)\)

OpenStudy (anonymous):

Solve for X

mathslover (mathslover):

ok . satellite is on right way

OpenStudy (anonymous):

then you can write your equation with only cosines you will get a quadratic to solve

OpenStudy (anonymous):

okay i will keep trying

OpenStudy (anonymous):

\[2\sin^2 (x )+ 3\cos(x) = 3\] \[2(1-\cos^2(x)) + 3\cos(x) = 3\] \[2(1-\cos^2(x)) + 3\cos(x)-3 = 0\] etc

OpenStudy (anonymous):

ohh okay! Thank you!